OpenStudy (anonymous):

I need help figuring this out. If f(x)=x^1\2-x and g(x)=2x^3-x^1/2-x, find f(x)-g(x). -2x^3-2x-2x^1/2 -2x^3-2x 2x^3-2x+2x^1\2 -2x^3+2x^1\2

OpenStudy (anonymous):

Do as it is asking \[f(x) - g(x)\] We know\[f(x) = x^{1/2}-x\] and \[g(x) = 2x^3-x^{1/2}-x\] So you must subtract g(x) from f(x) using those 2 equations.

OpenStudy (anonymous):

How do you know if it's either subtraction, addition, multiplication or division. That's what confuses me.

OpenStudy (anonymous):

It shows you in your original expression \[f(x) - g(x)\] It means to subtract g(x) from f(x)

OpenStudy (anonymous):

Okay... I just put it in the calculator and I got 2x^3-2x^1/2 but that's not one of the choices.

OpenStudy (anonymous):

You were close,

OpenStudy (anonymous):

You forget that you are subtracting so the first term is negative -2x^3

OpenStudy (anonymous):

and the second turns positive

OpenStudy (anonymous):

\[f(x) - g(x)\] \[f(x) = x^{1/2} - x\]\[g(x) = 2x^3-x^{1/2}- x\] So \[x^{1/2}- x - (2x^3 - x^{1/2} - x)\] You can see that they are all being subtracted so you get \[x^{1/2} - x -2x^3 + x^{1/2}+x\]

OpenStudy (anonymous):